OpenStudy (anonymous):

An oil company distributes oil in a metal can shaped like a cylinder that has an actual radius of 5.1cm and a height of 15.1cm. A worker incorrectly measured the radius as 5cm and the height as 15cm. Determine the relative error in calculating the surface area, to the nearest thousandth.

OpenStudy (anonymous):

What have you tried so far?

OpenStudy (agreene):

So, Surface Area of a Cylinder is: \[2\pi r^{2} + 2\pi rh\] where r is your radius and h is the height. From there we can plug in both sets of information. "Actual" measurements will lead you to: 647.294 "False" measurements will lead you to: 628.319 The error would be the difference, so: 647.294-628.319= 18.97 cm.

OpenStudy (anonymous):

That's the absolute error. They want the relative error.